OpenStudy (anonymous):

average velocity between t=2 & t=4? s(t) =t^3 -3t^2-9t+5

OpenStudy (anonymous):

(s(4)-s(2) )/ 2 = -16m/s? this correct?

myininaya (myininaya):

yes you do (s(4)-s(2))/(4-2) gj

OpenStudy (anonymous):

why would i not input my numbers in a v(t) formula tho?

OpenStudy (anonymous):

if the arithmetic is right, yes. it is \[\frac{s(4)-s(2)}{4-2}\]

myininaya (myininaya):

average velocity=slope of secant line velocity=slope of a tangent line two different things

OpenStudy (anonymous):

average mean 'average' \[\frac{distance traveled}{time elapsed}\]

OpenStudy (anonymous):

hello my myininaya!

myininaya (myininaya):

hey

OpenStudy (anonymous):

time for judge judy, see you soon i hope

myininaya (myininaya):

lol ok